# Entropy-Conservative numerical fluxes for compressible Euler equations with thermally perfect gas models

### v1

@@ 2.2 Comparison with existing schemes
$\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}+\overline{p},\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[e_{\mathrm{ref}}+c_{0}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{R}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}+\dfrac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{\left(p,u\right)}}.$ (11)
@@
This set of fluxes can be compared to the EC fluxes derived by Gouasmi et al. [19], who analysed the same problem with a polynomial dependence for $c_{v}(T)$ by using the theory by Tadmor [11, 33]. In the case of single-component gas, their formula reads, in our notation,
@@
$\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\dfrac{R\,\overline{\rho}\,}{\overline{1/T}\,}},\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[e_{\mathrm{ref}}+c_{0}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{R}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}+\dfrac{u_{i}u_{i+1}}{2}\right]+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{u}\,\dfrac{R\,\overline{\rho}\,}{\overline{1/T}\,}}.$ (12)
@@
By comparing Eqs. (11) and (12) one observes that the two sets of fluxes are almost coincident, except for the pressure terms in the momentum and total-energy equations, highlighted in red in Eq. (12). While in the present formulation (Eq. (11)) the pressure flux in the momentum equation is expressed through the arithmetic mean $\overline{p}\,$, in the formulation by Gouasmi et al. (Eq. (12)) it is expressed by using the complex average $\widehat{p}=R\,\overline{\rho}\,/\overline{1/T}\,$, which depends on density and temperature. Moreover, the pressure term in the total-energy equation is expressed as $\overline{\overline{(p,u)}}\,$ in Eq. (11) and as $\widehat{p}\,\,\overline{u}\,$ in Eq. (12). It is well known that the discretization of the pressure term in the momentum equation plays an important role in the global discrete preservation of kinetic energy. In fact, in [34], which deals with KEP split forms for calorically perfect gases, it is shown that even in the case in which the convective flux in the momentum equation is written in the KEP form as $\mathcal{F}^{\text{c}}_{\rho u}=\mathcal{F}_{\rho}\overline{u}\,$, the discretization of the pressure term can spoil the correct balance of induced kinetic energy causing errors in the pressure work contribution. This is especially true when, as in the case of the Chandrashekar scheme [14] to which Eq. (12) reduces for ideal gases, the pressure average is rendered as a combination of averages of density and temperature. The conclusions of the analysis conducted in [34] are that, in the case of calorically perfect gases, the simple arithmetic average of pressure in the momentum flux is the choice yielding the best results for the case of inviscid Taylor-Green vortices. In Sec. 4.2 we show that this result also applies in the case of thermally perfect gases. Similarly, the treatment of the pressure term in the total energy equation is known to be an element potentially affecting certain statistical quantities related to the fluctuating motion in homogeneous isotropic turbulence. In [35] and [30], as an example, it is shown that r.m.s of density and temperature fluctuations in homogeneous isotropic turbulence simulations can erroneously grow without bound if the pressure term in the total energy equation is different from $\overline{\overline{(p,u)}}\,$. In Sec. 4.2 we show that this is the case also for thermally perfect gases, demonstrating the importance of the set of fluxes here derived as a relevant improvement on existing formulations.
@@ 3.1 Numerical fluxes for hypersonic flows with the RRHO model
$c_{v}(T)=aR+R\dfrac{\frac{\theta^{2}}{T^{2}}e^{\frac{\theta}{T}}}{\left(e^{\frac{\theta}{T}}-1\right)^{2}},\qquad s=s_{\mathrm{ref}}+R\left(a\log T+\dfrac{\theta}{T}+\dfrac{\frac{\theta}{T}}{e^{\frac{\theta}{T}}-1}-\log(e^{\frac{\theta}{T}}-1)-\log\rho\right)+\text{const.},$. ,
@@ 3.2 Multicomponent flows
In the case of a gas composed of $N_{s}$ non-reacting species, the usual standard approach considers the mixture as a single compressible fluid, with density, momentum and total energy (per unit volume) given by the sum of the individual components, where the density of each species evolves according to a separate continuity equation. In this case, the compressible Euler equations are still in the form of Eq. (1) where $\mathbf{u}=\left(\rho_{1},\rho_{2},\ldots,\rho_{N},\rho u,\rho E\right)^{\text{T}}$ and $\mathbf{f}=\left(\rho_{1}u,\rho_{2}u,\ldots,\rho_{N_{s}}u,\rho uu+p,u(\rho E+p)\right)^{\text{T}}$, where $\rho_{k}=\rho Y_{k}$ is the density of the $k$-th species, $Y_{k}$ being the mass fraction and $\rho$ the density of the mixture. From the relation $\sum_{k}Y_{k}=1$ it follows $\sum_{k}\rho_{k}=\sum_{k}\rho Y_{k}=\rho$. Internal energy and entropy of the mixture are defined as $\rho e=\sum_{k}\rho_{k}e_{k}$ and $\rho s=\sum_{k}\rho_{k}s_{k}$, which also implies $e=\sum_{k}Y_{k}e_{k}$ and $s=\sum_{k}Y_{k}s_{k}$. Individual equations of state $p_{k}=\rho_{k}R_{k}T$, where $R_{k}$ is the gas constant of the $k$-th species, and Dalton’s law $\sum_{k}p_{k}=p$ are assumed. The Gibbs’ relation holds for each species: $d\left(\rho_{k}e_{k}\right)=Td\left(\rho_{k}s_{k}\right)+g_{k}d\rho_{k}$, and following the steps reported in [26], which lead to Eq. (3), one obtains the analogous sufficient conditions for entropy conservation:
@@
with $g_{k}=\omega_{k}(T)+R_{k}T\log{\rho_{k}}$ as in Eq. (4), where internal energy and entropy for each thermally perfect component are given by $e_{k}=\int_{T_{\mathrm{ref}}}^{T}c_{v\,k}(T^{\prime})\,\mathrm{d}{T^{\prime}}+e_{k\,\text{ref}}$ and $s_{k}=\int_{T_{\mathrm{ref}}}^{T}\frac{c_{v\,k}(T^{\prime})}{T^{\prime}}\,\mathrm{d}{T^{\prime}}-R_{k}\log{(\rho_{k}/\rho_{k\,\mathrm{ref}}})+s_{k\,\mathrm{ref}}$.
In the case of polynomial approximation for $c_{v\,k}(T)=\sum_{m=-L}^{R}c_{m}^{k}{T}^{m}$ the final set of fluxes reads
$\mathcal{F}_{\rho_{k}}=\overline{\rho_{k}}^{\text{log}}\overline{u}\,\qquad\!\mathcal{F}_{\rho u}=\sum_{k=1}^{N_{s}}\mathcal{F}_{\rho_{k}}\,\overline{u}+\overline{p},\qquad\!\mathcal{F}_{\rho E}=\sum_{k=1}^{N_{s}}\mathcal{F}_{\rho_{k}}\left[e_{k\,\mathrm{ref}}+c_{0}^{k}\dfrac{1}{\overline{1/T}^{\text{log}}}\!+\!\!\sum_{\begin{subarray}{c}m=1\end{subarray}}^{R}\frac{c_{m}^{k}}{m+1}\overline{\overline{T}}^{m}\!\!+\dfrac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{\left(p,u\right)}}.$ (17)
For calorically perfect gases, they reduce to
$\mathcal{F}_{\rho_{k}}=\overline{\rho_{k}}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\sum_{k=1}^{N_{s}}\mathcal{F}_{\rho_{k}}\,\overline{u}+\overline{p},\qquad\mathcal{F}_{\rho E}=\sum_{k=1}^{N_{s}}\mathcal{F}_{\rho_{k}}\left[e_{k\,\mathrm{ref}}+\left(\overline{1/e_{k}}^{\text{log}}\right)^{-1}+\dfrac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{\left(p,u\right)}},$ (18)
which is the extension of the Ranocha’s flux to multicomponent flows of ideal gases.
@@ 4 Numerical results
Numerical tests are presented to assess the EC properties and the robustness of the proposed formulation; all simulations have been carried out using the dimensionless form of the equations. Time integration has been performed with a standard fourth-order explicit Runge-Kutta method, with a sufficiently small CFL number to limit time-integration errors. For the thermodynamic framework, we have adopted the experimental fit for $c_{v}(T)$ given by the five-coefficient approximation reported in [32], where the only non-zero coefficients are $c_{-2},\,c_{0},\,c_{1},\,c_{2},\,c_{3}$. In the following results, the specific gas chosen is high-enthalpy methane, whose heat capacities strongly depend on temperature [37, 32]. Moreover, the non-polarity of such molecule allows Van der Waals forces to be neglected, especially at high temperatures where its modelling as a thermally perfect gas is preferable. Regarding the spatial discretization, we use the novel formulation (9)–(10) (together with the numerical flux for momentum given in Eq. (2)), here termed AEC-TP(N), with a sufficiently high value of $N$; in all our tests we checked that the value of $N$ guaranteeing the convergence to the EC scheme in Eq. (8) is typically not greater than $5$. The novel discretization is compared to the one proposed by Gouasmi et al., reported in Eq. (12) and generalized to include also negative exponents in the expression of the specific heat capacities. Additionally, results are provided for other widely used discretizations, such as the KEEP formulation [38], whose fluxes are
$\mathcal{F}_{\rho}=\overline{\rho}\,\overline{u}\,,\qquad\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}\,+\overline{p}\,,\qquad\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[\overline{e}\,+\frac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{(p,u)}}\,,$
the flux proposed by Ranocha [13]:
$\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,,\qquad\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}\,+\overline{p}\,,\qquad\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\,\left[\left(\overline{1/e}^{\text{log}}\right)^{-1}+\frac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{(p,u)}}\,$
@@
$\mathcal{F}_{\rho}=\overline{\rho}\,\overline{u}\,,\qquad\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}\,+\overline{p}\,,\qquad\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[\overline{e}\,+\overline{\frac{u^{2}}{2}}\,\right]+\mathcal{F}_{\rho}\overline{\left(\frac{p}{\rho}\right)}\,.$
In order to better evaluate the effect that the discretizations have on the evolution of quantities such as entropy and kinetic energy, we define a normalized difference of the integral value with respect to its value at time zero. For a generic variable $\phi$, that is

### v2

@@ 2.3 Comparison with existing schemes
$\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}+\overline{p},\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[\varepsilon_{\mathrm{ref}}+c_{0}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{r}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}+\dfrac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{\left(p,u\right)}}.$ (13)
@@
This set of fluxes can be compared to the EC fluxes derived by Gouasmi et al. [36], who analysed the same problem with a polynomial dependence for $c_{v}(T)$ by using the theory by Tadmor [21, 29]. In the case of single-component gas, their formula reads, in our notation [^1] It should be noted that, in the original formulation by Gouasmi et al. [36], $e_{\mathrm{ref}}$ was used in place of $\varepsilon_{\mathrm{ref}}$ since it was assumed that $T_{\mathrm{ref}}$ would be equal to zero.,
@@
$\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\dfrac{R\,\overline{\rho}\,}{\overline{1/T}\,}},\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[\varepsilon_{\mathrm{ref}}+c_{0}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{r}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}+\dfrac{u_{i}u_{i+1}}{2}\right]+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{u}\,\dfrac{R\,\overline{\rho}\,}{\overline{1/T}\,}}.$ (14)
@@
By comparing Eqs. (13) and (14) one observes that the two sets of fluxes are almost coincident, except for the pressure terms in the momentum and total-energy equations, highlighted in red in Eq. (14). While in the present formulation (Eq. (13)) the pressure flux in the momentum equation is expressed through the arithmetic mean $\overline{p}\,$, in the formulation by Gouasmi et al. (Eq. (14)) it is expressed by using the complex average $\widehat{p}=R\,\overline{\rho}\,/\overline{1/T}\,$, which depends on density and temperature. Moreover, the pressure term in the total-energy equation is expressed as $\overline{\overline{(p,u)}}\,$ in Eq. (13) and as $\widehat{p}\,\,\overline{u}\,$ in Eq. (14). It is well known that the discretization of the pressure term in the momentum equation plays an important role in the global discrete preservation of kinetic energy. In fact, in [18], which deals with KEP split forms for calorically perfect gases, it is shown that even in the case in which the convective flux in the momentum equation is written in the KEP form as $\mathcal{F}^{\text{c}}_{\rho u}=\mathcal{F}_{\rho}\overline{u}\,$, the discretization of the pressure term can spoil the correct balance of induced kinetic energy causing errors in the pressure work contribution. This is especially true when, as in the case of the Chandrashekar scheme [24] to which Eq. (14) reduces for ideal gases, the pressure average is rendered as a combination of averages of density and temperature. The conclusions of the analysis conducted in [18] are that, in the case of calorically perfect gases, the simple arithmetic average of pressure in the momentum flux is the choice yielding the best results for the case of inviscid Taylor–Green vortices. In Sec. 4.2 we show that this result also applies in the case of thermally perfect gases. Similarly, the treatment of the pressure term in the total energy equation is known to be an element potentially affecting certain statistical quantities related to the fluctuating motion in homogeneous isotropic turbulence. In [17] and [12], as an example, it is shown that r.m.s of density and temperature fluctuations in homogeneous isotropic turbulence simulations can erroneously grow without bound if the pressure term in the total energy equation is different from $\overline{\overline{(p,u)}}\,$. In Sec. 4.2 we show that this is the case also for thermally perfect gases, demonstrating the importance of the set of fluxes here derived as a relevant improvement on existing formulations.
@@ 3.1 Numerical fluxes for hypersonic flows with the RRHO model
$c_{v}(T)=aR+R\dfrac{\frac{\theta^{2}}{T^{2}}e^{\frac{\theta}{T}}}{\left(e^{\frac{\theta}{T}}-1\right)^{2}},\qquad s=s_{\mathrm{ref}}+R\left(a\log T+\dfrac{\theta}{T}+\dfrac{\frac{\theta}{T}}{e^{\frac{\theta}{T}}-1}-\log(e^{\frac{\theta}{T}}-1)-\log\rho\right)+\text{const.},$
@@ 3.2 Multicomponent flows
In the case of a gas composed of $N_{s}$ nonreacting species, the usual standard approach considers the mixture as a single compressible fluid, with density, momentum and total energy (per unit volume) given by the sum of the individual components, where the density of each species evolves according to a separate continuity equation. In this case, the compressible Euler equations are still in the form of Eq. (1) where $\mathbf{u}=\left(\rho_{1},\rho_{2},\ldots,\rho_{N},\rho u,\rho E\right)^{\text{T}}$ and $\mathbf{f}=\left(\rho_{1}u,\rho_{2}u,\ldots,\rho_{N_{s}}u,\rho uu+p,u(\rho E+p)\right)^{\text{T}}$, where $\rho_{k}=\rho Y_{k}$ is the density of the $k$-th species, $Y_{k}$ being the mass fraction and $\rho$ the density of the mixture. From the relation $\sum_{k}Y_{k}=1$ it follows $\sum_{k}\rho_{k}=\sum_{k}\rho Y_{k}=\rho$. Internal energy and entropy of the mixture are defined as $\rho e=\sum_{k}\rho_{k}e_{k}$ and $\rho s=\sum_{k}\rho_{k}s_{k}$, which also implies $e=\sum_{k}Y_{k}e_{k}$ and $s=\sum_{k}Y_{k}s_{k}$. Individual equations of state $p_{k}=\rho_{k}R_{k}T$, where $R_{k}$ is the gas constant of the $k$-th species, and Dalton’s law $\sum_{k}p_{k}=p$ are assumed. The Gibbs’ relation, which holds for each species: $d\left(\rho_{k}e_{k}\right)=Td\left(\rho_{k}s_{k}\right)+g_{k}d\rho_{k}$, can be used to relate the convective and pressure terms in the balance equations for mass, internal energy and entropy for each component of the mixture, as it was done for a single compound in [39]. By imposing a locally conservative formulation for mass and internal energy, and a straightforward discretization of the pressure term, one obtains the analogous sufficient conditions for entropy conservation of each species:
@@
with $g_{k}=\omega_{k}(T)+R_{k}T\log{\rho_{k}}$ as in Eq. (4), where internal energy and entropy for each thermally perfect component are given by $e_{k}=\int_{T_{\mathrm{ref}}}^{T}c_{v\,k}(T^{\prime})\,\mathrm{d}{T^{\prime}}+e_{k\,\text{ref}}$ and $s_{k}=\int_{T_{\mathrm{ref}}}^{T}\frac{c_{v\,k}(T^{\prime})}{T^{\prime}}\,\mathrm{d}{T^{\prime}}-R_{k}\log{(\rho_{k}/\rho_{k\,\mathrm{ref}}})+s_{k\,\mathrm{ref}}$. Eq. (18) is the general sufficient condition for entropy preservation of each species in the mixture.
In the case of polynomial approximation for $c_{v\,k}(T)=\sum_{m=0}^{r}c_{m}^{k}{T}^{m}$ the final set of fluxes reads
$\mathcal{F}_{\rho_{k}}=\overline{\rho_{k}}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\sum_{k=1}^{N_{s}}\mathcal{F}_{\rho_{k}}\,\overline{u}+\overline{p},\qquad\mathcal{F}_{\rho E}=\sum_{k=1}^{N_{s}}\mathcal{F}_{\rho_{k}}\left[\varepsilon_{k\,\mathrm{ref}}+c_{0}^{k}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{r}\frac{c_{m}^{k}}{m+1}\overline{\overline{T}}^{m}+\dfrac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{\left(p,u\right)}}.$ (19)
which can be compared to Eq. (44) of [36], where the same considerations on the discretization of the pressure terms still hold in this case.
For calorically perfect gases, the fluxes reduce to
$\mathcal{F}_{\rho_{k}}=\overline{\rho_{k}}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\sum_{k=1}^{N_{s}}\mathcal{F}_{\rho_{k}}\,\overline{u}+\overline{p},\qquad\mathcal{F}_{\rho E}=\sum_{k=1}^{N_{s}}\mathcal{F}_{\rho_{k}}\left[\varepsilon_{k\,\mathrm{ref}}+\left(\overline{1/e_{k}}^{\text{log}}\right)^{-1}+\dfrac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{\left(p,u\right)}},$ (20)
which is the extension of the Ranocha’s flux to multicomponent flows of ideal gases. To the best of the authors’ knowledge, this extension has not been mentioned in the open literature up to now.
@@ 4 Numerical results
Numerical tests are presented to assess the EC properties and the robustness of the proposed formulation; all simulations have been carried out using the dimensionless form of the equations. Time integration has been performed with a standard fourth-order explicit Runge–Kutta method, with a sufficiently small CFL number to limit time-integration errors. For the thermodynamic framework, we have adopted the experimental fit for $c_{v}(T)$ given by the five-coefficient approximation reported in [42], where the only nonzero coefficients are $c_{-2},\,c_{0},\,c_{1},\,c_{2},\,c_{3}$. In the following results, the specific gas chosen is high-enthalpy methane, whose heat capacities strongly depend on temperature [42, 43]. Moreover, the nonpolarity of such molecule allows van der Waals forces to be neglected, especially at high temperatures where its modelling as a thermally perfect gas is preferable. Regarding the spatial discretization, we use the novel formulation (11)–(12) (together with the numerical flux for momentum given in Eq. (2)), here termed AEC-TP(N), with a sufficiently high value of $N$; in all our tests we checked that the value of $N$ guaranteeing the convergence to the EC scheme in Eq. (10) is typically not greater than $5$. The novel discretization is compared to the one proposed by Gouasmi et al., reported in Eq. (14) and generalized to include also negative exponents in the expression of the specific heat capacities. Additionally, results are provided for other widely used discretizations, such as the formulation proposed by Ranocha [23]
$\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,,\qquad\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}\,+\overline{p}\,,\qquad\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\,\left[\left(\overline{1/e}^{\text{log}}\right)^{-1}+\frac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{(p,u)}}\,$, (21)
the KEEP formulation proposed by Kuya et al. [45], whose fluxes are
$\mathcal{F}_{\rho}=\overline{\rho}\,\overline{u}\,,\qquad\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}\,+\overline{p}\,,\qquad\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[\overline{e}\,+\frac{u_{i}u_{i+1}}{2}\right]+\overline{\overline{(p,u)}}\,$, (22)
@@
$\mathcal{F}_{\rho}=\overline{\rho}\,\overline{u}\,,\qquad\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}\,+\overline{p}\,,\qquad\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[\overline{e}\,+\overline{\frac{u^{2}}{2}}\,\right]+\mathcal{F}_{\rho}\overline{\left(\frac{p}{\rho}\right)}\,.$ (23)
The Ranocha’s flux in Eq. (21) is exactly EC and PEP for calorically perfect gases, and is also KEP. The KEEP and JP fluxes in Eq. (22) and (23) are KEP, but not EC, although the KEEP formulation has noticeable entropy conservation properties for calorically perfect gases [17, 27, 26]. In order to better evaluate the effect that the discretizations have on the evolution of quantities such as entropy and kinetic energy, we define a normalized difference of the integral value with respect to its value at time zero. For a generic variable $\phi$, that is